Can you please help me solve the following equation: L= 5more than breath P= 30 what is the length and breath of the rectangular pool ?

Answer by nerdybill(7384)

Can you please help me solve the following equation: L= 5more than breath P= 30 what is the length and breath of the rectangular pool ?
since
2(length + breath) = perimeter
Let B = breath
then
B+5 = length
.
so we have:
2(length + breath) = perimeter
2(B+5 + B) = 30
2(2B+5) = 30
2B+5 = 15
2B = 10
B = 5 (breath of pool)
.
Length of pool:
B+5 = 5+5 = 10